# OmniAD: View-Invariant Unsupervised Scene Pixelwise Anomaly Detection with Adaptive View Synthesis

### v2

@@ 5.5 Qualitative Evaluation
OmniAD-R 0.999 0.993 0.999 0.999 0.988 0.983 0.999 0.999 0.924 0.892 0.933 0.928
OmniAD-$A^{i}$ 0.999 0.999 0.998 0.999 0.983 0.988 0.997 0.997 0.948 0.946 0.966 0.963
RD 0.999 0.999 0.999 0.999 0.981 0.982 0.998 0.998 0.936 0.929 0.951 0.956
CFLOW 0.996 0.996 0.999 0.999 0.962 0.967 0.993 0.990 0.945 0.906 0.947 0.953
FastFlow 0.999 0.998 0.999 0.999 0.868 0.896 0.965 0.966 0.862 0.914 0.934 0.932
CFA 0.998 0.998 0.999 0.994 0.977 0.976 0.986 0.979 0.916 0.927 0.934 0.942
DRÆM 0.819 0.359 0.514 0.524 0.720 0.571 0.684 0.796 0.601 0.496 0.559 0.519
@@ 6 Limitations and Future Work
OmniAD and the proposed ToyCity dataset address unsupervised Scene Anomaly Detection in multi-view scenarios, yet several limitations remain. First, each ToyCity image contains at most one anomaly at a time, even though multiple anomalies were added and removed from each scene. In contrast, real-world settings such as disaster assessments or industrial inspections may involve multiple anomalies within a single capture. Note that ToyCity does include partial occlusions, and OmniAD itself does not strictly limit detection to a single anomaly region. Second, although ToyCity was captured with real cameras at varying angles and lighting conditions, its tabletop-scale scenes do not fully encompass the complexities of large scale infrastructure.
@@
    Investigate faster 3D reconstruction, localization, and noval view synthesis methods.
  • 4.
    Identifying the underlying failure modes in GAN/autoencoder based AD methods under significant viewpoint variation remains an open research question, which we leave for future work.
@@ 7 Conclusion
Our study introduced and addressed a specific anomaly detection (AD) problem we term Scene AD. Scene AD deals with pixel-level AD between a non-anomalous scene and an anomalous scene, where a misalignment between non-anomalous and anomalous views is expected, as is often the case in applications in the built environment. To challenge and measure the effectiveness of state-of-the-art anomaly detection for Scene AD, we introduced the ToyCity dataset. This new benchmark focuses on viewpoint variance in scenes of multiple objects and serves as a testbed for our new method OmniAD. We demonstrated through extensive comparisons that OmniAD outperforms other AD methods when applied to the ToyCity dataset as well as the object centric established MAD-Real, establishing a new state-of-the-art for this task. Particularly, we found that the incorporation of a student attention module increases the efficacy of reverse distillation for pixel-level AD by expanding the effective receptive filed (ERF). Our experiments also demonstrated the benefits of novel view synthesis and camera localization for dataset augmentation in enhancing the robustness and accuracy of OmniAD allowing for an increase in pixel level $F_{1}$ of up to 40% in many scenes. We hope that our work will catalyze further advancements, spurring the creation of advanced anomaly detection techniques capable of managing the complexities associated with varying viewpoints in 3D scenes.

### v3

@@ 5.4 Qualitative Evaluation
OmniAD w/o R 0.999 0.993 0.999 0.999 0.988 0.983 0.999 0.999 0.924 0.892 0.933 0.928
OmniAD w/o $A^{i}$ 0.999 0.999 0.998 0.999 0.983 0.988 0.997 0.997 0.948 0.946 0.966 0.963
RD [25] 0.999 0.999 0.999 0.999 0.981 0.982 0.998 0.998 0.936 0.929 0.951 0.956
CFLOW-AD [21] 0.996 0.996 0.999 0.999 0.962 0.967 0.993 0.990 0.945 0.906 0.947 0.953
FastFlow [22] 0.999 0.998 0.999 0.999 0.868 0.896 0.965 0.966 0.862 0.914 0.934 0.932
CFA [23] 0.998 0.998 0.999 0.994 0.977 0.976 0.986 0.979 0.916 0.927 0.934 0.942
DRÆM [24] 0.819 0.359 0.514 0.524 0.720 0.571 0.684 0.796 0.601 0.496 0.559 0.519
Table 11: Comparison of pixel wise AUROC/$F_{1}$ metrics on the MVTec-AD[13] dataset.
Method OmniAD RD[25] UniAD[53] SimpleNet[51] AnomalyDINO[52]* MambaAD[34]
Objects Bottle 0.987/0.755 0.986/0.741 0.983/0.707 0.980/0.727 0.988/0.766
Cable 0.983/0.633 0.968/0.573 0.974/0.569 0.974/0.600 0.980/0.617
Capsule 0.984/0.509 0.989/0.499 0.980/0.401 0.989/0.491 0.986/0.475
Hazelnut 0.993/0.710 0.988/0.608 0.983/0.561 0.979/0.503 0.990/0.639
Metal Nut 0.987/0.865 0.969/0.795 0.944/0.688 0.988/0.866 0.971/0.806
Pill 0.983/0.703 0.976/0.667 0.952/0.519 0.985/0.726 0.973/0.660
Screw 0.993/0.479 0.995/0.472 0.984/0.236 0.993/0.388 0.995/0.517
Toothbrush 0.992/0.627 0.990/0.609 0.985/0.521 0.985/0.529 0.990/0.607
Transistor 0.967/0.632 0.906/0.549 0.988/0.585 0.968/0.621 0.968/0.675
Zipper 0.988/0.686 0.987/0.590 0.965/0.434 0.989/0.654 0.981/0.588
Textures Carpet 0.986/0.666 0.991/0.609 0.985/0.533 0.980/0.478 0.992/0.634
Grid 0.994/0.523 0.993/0.477 0.943/0.315 0.988/0.393 0.992/0.488
Leather 0.995/0.575 0.993/0.467 0.991/0.419 0.992/0.456 0.993/0.495
Tile 0.956/0.690 0.951/0.597 0.902/0.487 0.972/0.688 0.931/0.523
Wood 0.955/0.624 0.948/0.495 0.935/0.439 0.940/0.481 0.939/0.477
Mean 0.983/0.645 0.975/0.583 0.966/0.506 0.980/0.573 0.979/0.618 0.978/0.598
* Only mean was available from source.
5.5 Ablation Studies
We conducted ablation studies to validate both the proposed OmniAD architecture and the effectiveness of the data augmentation strategies.
5.5.1 OmniAD Architecture Ablation
To validate the architectural changes proposed in OmniAD and quantify the contribution of each key modification compared to the baseline Reverse Distillation [25], we conducted an ablation study on both OmniAD’s components and data augmentation effects. To evaluate the component wise effect of OmniAD we shall refer to the ablation configurations as:
  • 1.
    OmniAD: Full model (ResNeXt backbone + $A^{i}$).
  • 2.
    OmniAD w/o R: Original Reverse Distillation backbone + $A^{i}$ (isolates $A^{i}$ contribution without ResNeXt).
  • 3.
    OmniAD w/o $A^{i}$: ResNeXt backbone without $A^{i}$ (isolates ResNeXt contribution without $A^{i}$).
  • 4.
    RD: Baseline Reverse Distillation model (without $A^{i}$ and ResNeXt).
Results for these configurations are presented in Tables 5, 6, 7, 8, 9, and 10. We aggregate the effects of each component of OmniAD without augmentation in Table 12.
Table 12: Average Pixel $F_{1}$ Scores across ToyCity and MAD-Real[12] with no augmentation and percentage improvement over baseline RD.
Method $F_{1}$ Improvement over RD
OmniAD 0.398 32.5%
OmniAD w/o R 0.376 25.1%
OmniAD w/o $A^{i}$ 0.334 11.1%
RD[25] 0.300 0.0%
Contribution of Student Attention Modules ($A^{i}$): By comparing configurations with and without $A^{i}$ while keeping the backbone consistent (OmniAD w/o R vs. RD), we observe the impact of the student attention module. As shown in Table 12, the inclusion of $A^{i}$ leads to a 25.1% gain over RD. This supports our hypothesis that expanding the ERF via the student attention modules significantly enhances the model’s ability to localize anomalies.
Contribution of ResNeXt Backbone: The impact of the ResNeXt backbone is assessed by comparing OmniAD and RD configurations. As shown in Table 12, by comparing different backbones (OmniAD vs. OmniAD w/o $A^{i}$), using ResNeXt provides an 11.1% gain.
In summary, the architecture ablation study quantitatively demonstrates that both the ResNeXt backbone and the proposed student attention modules ($A^{i}$) contribute positively to the performance of OmniAD. The combination of these components in OmniAD leads to the superior performance observed in our experiments over Reverse Distillation.
5.5.2 NVS Augmentation Ablation
To determine the effectiveness of our Novel View Synthesis (NVS) data augmentation techniques, specifically Query-Aligned Non-anomalous Views (QANV) and Interpolated Non-anomalous Views (INV), we performed an ablation study on OmniAD’s performance. The underlying assumption was that augmenting the training data with synthesized non-anomalous views, either aligned with query images or interpolated between existing views, would enhance the model’s ability to generalize across different viewpoints and consequently improve anomaly detection. Table 13 presents a summary of OmniAD’s average pixel $F_{1}$ scores across all datasets under different augmentation conditions.
Table 13: Average $F_{1}$ score across ToyCity and MAD-Real[12] for different augmentation strategies.
Method Augmentation Strategy $F_{1}$ Improvement
OmniAD Both 0.493 64.33%
OmniAD QANV 0.493 64.33%
OmniAD INV 0.460 53.33%
OmniAD No Aug 0.398 32.67%
RD[25] No Aug 0.300 0%
The results in Table 13 clearly indicate that NVS-based augmentation markedly boosts OmniAD’s average performance. Both the QANV strategy and the combined use of QANV and INV (Both) achieved the highest average $F_{1}$ score of 0.493, which is approximately a 64.33% relative increase compared to Reverse Distillation without augmentation. The INV strategy, when used alone, also demonstrated a significant improvement, yielding an average $F_{1}$ score of 0.460.
The detailed results (Tables 5, 6, 9) reveal that the best performing augmentation (QANV, INV, or Both) can vary depending on the specific dataset and metric of interest. QANV frequently is shown to be the optimal in terms of $F_{1}$ for a specific dataset, while INV and ”Both” are frequently better in terms of AUROC. However, in real-world applications of Scene AD, selecting the empirically optimal augmentation strategy per dataset is not feasible due to the absence of ground truth labels. Fortunately, combining both QANV and INV offers a robust strategy that does not require dataset-specific optimization. The aggregated results across datasets shown in Table 13 shows that the combination of QANV and INV achieves the same high average $F_{1}$ score as QANV across all datasets. Thus, while combining QANV and INV may not be the most optimal for a specific dataset, the combination is optimal on average.
@@ 6 Limitations and Future Work
OmniAD and the proposed ToyCity dataset address unsupervised Scene Anomaly Detection in multi-view scenarios, yet several limitations remain. First, each ToyCity image contains at most one anomaly at a time, even though multiple anomalies were added and removed from each scene. In contrast, real-world settings such as disaster assessments or industrial inspections may involve multiple anomalies within a single capture. Note that ToyCity does include partial occlusions, and OmniAD itself does not strictly limit detection to a single anomaly region. Second, although ToyCity was captured with real cameras at varying angles and lighting conditions, its tabletop-scale scenes do not fully encompass the complexities of large-scale infrastructure.
@@
    Investigate faster 3D reconstruction, localization, and novel view synthesis methods.
@@ 7 Conclusion
This paper introduced and addressed the task of Scene Anomaly Detection (Scene AD), focusing on localizing anomalies within multi-object scenes captured through unlabeled images exhibiting significant, arbitrary viewpoint variations. We identified the limitations of existing datasets and methods, which typically assume fixed cameras or single-object focus. To bridge this gap, we presented ToyCity, the first multi-object, multi-view real-image dataset specifically designed to evaluate unsupervised AD methods for Scene AD. Furthermore, we demonstrated that established unsupervised AD techniques struggle with pixel-level localization on this challenging dataset. We proposed OmniAD, a novel anomaly detection architecture that refines the Reverse Distillation framework. By incorporating a more robust ResNeXt backbone and, critically, introducing student attention modules to expand the student decoder’s effective receptive field (ERF), OmniAD significantly enhances sensitivity to localized anomalies across varying perspectives. Complementing the architecture, we developed two augmentation strategies, Interpolated Non-anomalous Views (INV) and Query-Aligned Non-anomalous Views (QANV), which utilize a Neural Radiance Field for novel view synthesis (NVS) and hierarchical localization to generate targeted non-anomalous training views. Extensive experiments and ablation studies conducted on our ToyCity dataset and the real image subset of MAD validated our approach, showing that OmniAD with QANV and INV improves pixel-level anomaly localization ($F_{1}$ scores) by 64.33% over Reverse Distillation without QANV and INV. The results underscored the distinct contributions of the ERF expansion via student attention modules and the effectiveness of NVS-based augmentations. We hope our work provides a valuable initial benchmark for future research in view invariant anomaly detection for real-world scenes.